x(t) = (1 – t). [u(t) - ut - 1)] h(t) = t. [u(t+ 2) u(t - 2)]

SIGNALS AND CONVULTION QUESTION... PLEASE DONT TOUCH IF YOU ARE NOT SURE ABOUT SOLUTION...

First; draw x and h signals…

For ?(?) = ?(?) ∗ ℎ(?) convulsion Write the mathematical expressions of the ? (?) response.
Answer #1

Solution of this question is shown belowDATE SOL given xit) = (1-t). [uct) - uct-1)] > Graph of [uct) - ult-1)] 1 & Graph of (1-t) I 1 & Graph of (1-t) x (uct) - ultDATE Graph of Cart 2) -act-2)] & 12 (111 Graph of t O Graph of t. [ult+2) - ult-2)] hit) 2 2 2 git) 24* hlt) Formula - TILt =X(+-) DATE xlt-t) for tc-2 ovoslabing lylt) = 0 for -2 < t < 3 y(t) = Ex{1-t +c] at - 2 3 Yt) = full-t)e + e? Zac ght) = (1-2DATE No overlapping for t73 (010000 Y10) - So tc-2 y(t) = -2t <3 O t> 3 * ab 2020-6-6 02:16